The Metropolitan corporation has issued a total of 2,400,000 shares. The north side investment group owns 7.5% of those shares. How many shares does north side own?
Answer by stanbon(75887)

The Metropolitan corporation has issued a total of 2,400,000 shares. The north side investment group owns 7.5% of those shares. How many shares does north side own?
------
Ans: [(7.5)/100]*2,400,000 = 7.5*24000 = 180,000
Cheers,
Stan H.